OpenStudy (anonymous):

Henry and Irene working together can wash all the windows of their house in 3 h 28 min. Working alone, it takes Henry 3.5 h more than Irene to do the job. How long does it take each person working alone to wash all the windows?

OpenStudy (anonymous):

it can't be 4.5 hours for both of them because it says it takes Henry 3.5 hrs longer

OpenStudy (saifoo.khan):

\[\frac{1}{210+x}+\frac{1}{x}=\frac{1}{208}\]

OpenStudy (saifoo.khan):

Solve for x.

OpenStudy (anonymous):

so I find a common denominator: x(210+x)(208)?

OpenStudy (saifoo.khan):

No, there is an equal sign in between.

OpenStudy (saifoo.khan):

\[\frac{x+x(210+x)}{x(x+210)}=\frac{1}{208}\]

OpenStudy (anonymous):

simplified it would be 2x/x =1/208?

OpenStudy (saifoo.khan):

No!

OpenStudy (anonymous):

help...? I'm confused

OpenStudy (saifoo.khan):

\[\frac{x+210x + x^2}{x^2+210x} = \frac{1}{208}\]

OpenStudy (saifoo.khan):

Can u solve it now?

OpenStudy (anonymous):

I'm not doing this right I ended up with x/1=1/208